OpenStudy (anonymous):

Help! Please/Thanks :) Rob can clean an attic in 13 hours. Julia can clean the same attic in 16 hours. if they worked together how long would it take them? a) 6.48 hours b) 7.17 hours c) 7.76 hours d) 8.8 hours

OpenStudy (anonymous):

Use reciprocals when approaching these questions: ie. rate of cleaning is (1Attic)/(13hours) or (1/13)Attics/hour. You can combine their rates and invert for your answer. \[rate_{1}+rate_{2}=rate_{total}\] \[\frac{1}{13}+\frac{1}{16} = rate_{total}\] \[\frac{29}{208} = rate_{total} = \frac{Attic}{Hour}\] invert and you will have units of hours/attic and get your answer.

OpenStudy (anonymous):

you just saved my life, i got 7.17, thanks!

OpenStudy (anonymous):